A strain of peas has three green peas and one yellow in every four peas. If twelve peas are randomly selected, what is the probability that exactly 8 of the peas are green?
Answer #1

P(green), p = 3/4

q = 1 - 3/4 = 1/4

Sample size, n = 12

Binomial distribution: P(x) = nCx px qn-x

P(exactly 8 of the peas are green), P(8) = 12C8 x (3/4)8 x (1/4)4

= 0.1936